OpenStudy (anonymous):

A single die is rolled.find the probablity of rolling an even number or a number less then 5.

OpenStudy (anonymous):

In probability if it is 'or' you have to add all probility

OpenStudy (anonymous):

so what would you add?

OpenStudy (anonymous):

all the chances that it is even + all the chances that it is number lower than 5

OpenStudy (anonymous):

so 5+4+3+2+1?

OpenStudy (anonymous):

What are you chances of rolling an even number 3 out of 6 (2,4,6) so 3/6 What are you chances of rolling an number lower than 5 5/6

OpenStudy (anonymous):

you have six chances total and out of those you have three even numbers, so you make a fraction of the desired outcome divided by the total possible outcomes.

OpenStudy (m):

5/6?

OpenStudy (anonymous):

Remember that probilbilty is out of 1 so it can be bigger than 1

OpenStudy (anonymous):

I think Imranmeah91 meant can't

OpenStudy (anonymous):

Since what you are describing in all cases I think answer should be 1

OpenStudy (dumbcow):

P(even) = 3/6 = 1/2 P(less than 5) = 4/6 = 2/3 P(even and less than 5) = 2/6 = 1/3 P(even OR less 5) = 1/2 + 2/3 - 1/3 = 5/6

OpenStudy (anonymous):

less than 5 doesn't include 5, so 4 out of 6 = 4/6 = 2/3

OpenStudy (m):

even is included so includes 6

OpenStudy (anonymous):

but don't u have to reduce it?

OpenStudy (m):

no unless specifically stated

OpenStudy (anonymous):

even = 2, 4, 6, less than five = 4, 3, 2, 1 BUT WAIT! you already counted 2 and 4! so it is 1, 2, 3, 4, and 6 for 5 desired outcomes out of 6 possibilities = 5/6

OpenStudy (m):

dumbcow's answer is complete. more like proof lol.